OpenStudy (anonymous):

perform the indicated operation (9+7i)(9-7i) answer in the form a+bi

OpenStudy (anonymous):

\[(a+bi)(a-bi)=a^2+b^2\]

OpenStudy (anonymous):

81-63i+63i-49i^2

OpenStudy (anonymous):

\[9^2+7^2=81+49\]

OpenStudy (anonymous):

(9+7i)(9-7i) 81-63i+63i-49i^2 81-49i^2

OpenStudy (anonymous):

(9+7i)(9-7i) =81-63i+63i =81-49i^2 =130+0i ans....